a truck eneters the highway at 60mph. a car enters the highway 9 min. later at 65mph. How long until the car will pass the truck?
Answer by dabanfield(803)

a truck eneters the highway at 60mph. a car enters the highway 9 min. later at 65mph. How long until the car will pass the truck?
Let T be the time when the car passes the truck.
Using the fact that distance = rate * time we have:
The disatance traveled by the truck in the 9 minutes or 9/60 hours before the car starts is :
60*9/60 = 9 miles.
The car must travel the 9 miles plus the additional distance the truck travels after the car starts before they meet:
The distance truck travels before they meet is 9 + 60*T. In that same time T the car will travel 65*T Miles.
They meet when:
9 + 60*T = 65*T
Solve for T .